from a pointA (1,1),straight lines AL and AM are drawn at right angles to the pair of straight lines 3x2+7xy-2y2=0...then combined equation of AL and AM is............

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shivanshi

3x2+7xy-2y2=0

m1 + m2 = 7/2

 m1.m2 =-3/2

 

now let a line  a'(x-1)2 + 2h'(x-1)(y-1) + b'(y-1)2 =0   whose line is perpendicular to the these lines

 m3 = -1/m1

m4 = -1/m2

 

m3 + m4 = -(m1+m2)/m1m2  

 -2h'/b'         = 7/3

m3.m4  = -1/m1m1

   a'/b'     = -2/3

 

a'(x-1)2 + 2h'(x-1)(y-1) + b'(y-1)2 =0

or

a'/b' (x-1)2 + 2h'/b' (x-1)(y-1) + (y-1)2 =0

- 2/3  (x-1)2 -7/3 (x-1)(y-1) + (y-1)2 =0

or

   2 (x-1)2 +7 (x-1)(y-1) +3 (y-1)2 =0

 Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin